I've been trying to help my daughter with this but my brain if frozen.
you want to make a rectangular prism out of cardboard. the base is 3 feet by 4 feet. the height is 2 feet. you have $5 and cardboard costs $0.10 per square foot. Do you have enough money? If not how much more do you need?

Answer by stanbon(75887)

make a rectangular prism out of cardboard. the base is 3 feet by 4 feet. the height is 2 feet. you have $5 and cardboard costs $0.10 per square foot. Do you have enough money? If not how much more do you need?
----------
Surface area = 2[3*4+3*2+4*2] = 2*26 = 52 sq ft.
Cost = 0.10*52 = $5.20
-----
Ans: Not enough ; Need $0.20 or 20 cents
-----------------
Cheers,
Stan H.
---------------